laura is 4 years older than her bro.two years ago laura was three times as old as her brother how old are they now ?
Answer by Greenfinch(383)

Let Laura's age now be X,her brother is now X - 4
2 years ago X - 2 = 3(X - 6)
X - 2 = 3X - 18
X = 3X - 16
2X = 16
X = 8
To check Laura is now 8, so her brother is now 4
2 years ago, Laura was 6, brother was 2 which is 3 times.